OpenStudy (icalibear):

A class has 20 students. The mean score on a test for 19 of the students was 76. The 20th student earned a score of 96. What is the new mean score for the class?

OpenStudy (whpalmer4):

If the mean score for the first 19 students was 76, we can pretend that they all got 76, because 76 19 times has the same mean. Our new mean score would be \[\frac{\text{sum of all of the scores}}{\text{number of scores}} = \frac{19*76 + 96}{20}=\]